Suppose the level of output that creates full employment in the economy is 1800, using the multiplier, datermine the level of investment spending that will create full employment in this economy



1
Expert's answer
2020-11-17T07:18:52-0500

Aggregate expenditure"=C+I+G+(X-M)\newline1800=200+0.9(1800-0.3(1800)+600+G-0.1(1800-0.3(1800))\newline1800=200+0.9(1260)+600+G-0.1(1260)\newline1800=200+1134+600+G-1134\newline1800=800+G\newline G=1800-800\newline=1000"